OpenStudy (anonymous):

Solve (x^2+x-30)/(x-5)=11. Check your answer.

OpenStudy (anonymous):

First let's factor... \[x^2+x-30=(x-5)(x+6)\] \[(x-5)(x+6)/(x-5)=11\]\[x+6=11\]\[x=5\]

OpenStudy (anonymous):

I got it now! Thanks!